A fire cracker was launched in the air at an angle of 750 with an initial velocity around 18.5m/s.

a. How high can the fire cracker reach?
1
Expert's answer
2020-12-14T07:17:39-0500

The maximum height is


"H=\frac{v^2\text{ sin}^2\theta}{2g}=16.3\text{ m}."